OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and that passes through the given point. y=9/5x-5;(10,5)

OpenStudy (campbell_st):

your gradient will be m = 9/5 so the line is in the form y = 9/5 x + b substitute x = 10 and y = 5 , from your point, into the equation above to find b. then you will have the equation of the line.

OpenStudy (anonymous):

y = 9/5x – 13

OpenStudy (campbell_st):

correct